OpenStudy (anonymous):

How many distinct real solutions exist for x, given log(20x) + log(50x) - 4log(x - 1) = 3? Please explain ho you figured it out!

OpenStudy (anonymous):

i'm assuming all this is log base 10?

OpenStudy (anonymous):

Yes. They are all common logarithms.

OpenStudy (anonymous):

lol... look at your question... I'm not a ho

OpenStudy (anonymous):

lol Sorry. typo. *how :)

OpenStudy (anonymous):

\(\large log(20x)+log(50x)-4log(x-1)=3 \) \(\large log(\frac{(20x)(50x)}{(x-1)^4})=3 \) \(\large 10^3=\frac{(20x)(50x)}{(x-1)^4} \) \(\large 10^3=\frac{(1000x^2)}{(x-1)^4} \) \(\large 1=\frac{(x^2)}{(x-1)^4} \) \(\large (x-1)^4=x^2 \) \(\large x^4-4x^3+6x^2-4x+1=x^2 \) \(\large x^4-4x^3+5x^2-4x+1=0 \)

OpenStudy (anonymous):

that last line can be factored as: \[\large (x^2-3x+1)(x^2-x+1)=0 \] i guess you can use the quadratic formula for each factor...

OpenStudy (anonymous):

Oh, wow! But how did you manage to factor it into the two trinomials?

OpenStudy (anonymous):

uh.. mr. wolfram...:)

OpenStudy (anonymous):

sorry... i can't factorize into trinomials...

OpenStudy (anonymous):

lol. I'm sure I'll figure it out eventually...Thanks though! By the way, how do you get a line underneath the fraction when writing an equation? I can get the forward slash, but not like a fraction.

OpenStudy (anonymous):

in this box just use "\[ \ and end it with \]" without the quotes. right click this and choose show math as : \[\large \frac{\sqrt{3}}{2}\]

OpenStudy (anonymous):

tex commands....

OpenStudy (anonymous):

or you can just use the equation button , right next to the draw button at the lower left of this box.

OpenStudy (anonymous):

ALright. Thanks! :)

OpenStudy (anonymous):

yw...:)